1) Assume ¥102.36 equal $1. Also assume that SKr6.5103 equal $1. How many Japanese yen can you acquire in exchange for 5,000 Swedish kronor?

A ¥318

B ¥261

C ¥49,520

D ¥78,614

2) Which one of the following statements related to the capital asset pricing model approach to equity valuation is correct? Assume the firm uses debt in its capital structure.

A The model generally produces the same cost of equity as the dividend growth model.

B The model is dependent upon a reliable estimate of the market risk premium.

C This approach generally produces a cost of equity that equals the firm's overall cost of capital.

D The model applies only to non-dividend-paying firms.
Answer #1

Answer: D Yen 78614

1 $ Yen 102.36
1$ SKr 6.5103
Yen/ SKr = Yen/ $ * $/SKr
Yen/ SKr = 102.36 * (1/6.5103)
1 SKr = Yen (102.36/6.5103)
5000 SKr = Yen (5000*(102.36/6.5103))
5000 SKr = Yen 78614

Answer (B)

The model is dependent upon a reliable estimate of the market risk premium.

capital asset pricing model approach to equity valuation

Requited Return = Risk Free Rate+ Beta*Market Risk Premium